Calculate the pH of a 500.0 mL solution that is 0.30 M NH3 and 0.50 M NH4Br. For NH3, Kb = 1.76 × 10-5 .
Answer #1

рь - - Дое % -lo (1.76 х 10) - 4,2ѕч рои - ее о р ) = 4. 154 + goo | O, SO О, 20 ) - 4.97 от 14 - орон = 14 - Ч.97 = 9.0 2 4

If you have any questions please comment

If you satisfied with the solution please rate it thanks